how do I figure this out using excel? Problem 10-02 Algo (What If Analysis) Question 5 of 22 Check My Work eBook Cox Electric makes electronic components and has e
Answer #1

Fixed Cost, F = 7000

Variable cost, v = 0.15 + 0.10 = 0.25

Selling Price, s = 0.65

Let x be the production volume

Profit = Total Revenue - Cost Price

= s*x - ( F + v*x)

= 0.65x - 7000 - 0.25x

= 0.40x - 7000

Now puttinx x from 5000 to 50000 with 5000 difference

X ( Production Volume ) Profit   
5000 -5000
10000 -3000
15000 -1000
20000 1000
25000 3000
30000 5000
35000 7000
40000 9000
45000 11000
50000 13000

Going from negative for x = 15000 to x = 20000

a) 15000 to 20000

b) Using Goal seek

Exact Break Even Point = 17,500

b) 17500

If any doubt please feel free to ask through the comments section and if no doubt please give a like. Thank You